### v2

@@ 4 Weakly Shill-Proof and Strategy-Proof Auctions
For all $\varepsilon>0$, there exists a value distribution $F$ and screen level $v^{Y}$ such that $Q^{AS,Y}(F)/Q^{E}(F)<\varepsilon$ and the ascending, screening auction with screening level $v^{Y}$ is weakly shill-proof, ex-post incentive compatible, and optimal.
The ascending, screening auction is orderly and optimal by construction; and is ex-post incentive compatiblebecause the English auction phase and the second-price phase both induce the same (ex-post incentive compatible) allocation and transfer rule. The larger $v^{Y}$ is, the less that can be extracted in expectation from shill bidding and the more likely it is that a shill bidder will win the item if she shill bids. We provide a sufficient minimum bound on $v^{Y}$ based on a few moments of a distribution (not its number of atoms), such that for distributions with “thin-enough” right tails—in particular monotone hazard rate distributions— the ascending, screening auction is weakly shill-proof (Lemma C.3). We can then construct a sequence of distributions with increasing numbers of atoms and constant $v^{Y}$ to complete the proof.
5 Single-Action Auctions
Theorem 4.5 shows, by “compressing” an English auction, that there exists a weakly shill-proof and ex-post incentive compatible auction in which, for some value distributions, bidders take far fewer actions than in an English auction. We now show that such compression has its limits, and more generally that there is no weakly shill-proof, ex-post incentive compatible, and optimal auction in which each bidder takes a single action. This impossibility result holds even after relaxing the assumptions that $F$ is regular and after weakening our concept of ex-post incentive compatiblility.
5.1 Set-Up
Let us begin by defining a single-action auction. An auction is considered single-action when each bidder takes precisely one action in the auction (under all possible histories). More formally, for any $h_{N}\in Z$, let $h_{\emptyset}\prec h_{1}\prec\ldots h_{N-1}\prec h_{N}$ be the sequence of histories to reach $h_{N}$. Then, for all $i\in B$, there exists a unique $n\leq N$ such that $i=P(h_{n})$. Without loss, we label the bidders $1,\ldots,N$, in the order that they move and label the action taken by bidder $i$ as $a_{i}$. [^31] Note that this labeling need not be the same for different histories as the bidder ordering can be endogenous to actions taken.
For exposition purposes, instead of tracking the information set $\mathcal{I}_{i}$ of a bidder $i\in R$, we assume that the information $i$ has when taking an action is a signal $s_{i}\in\mathcal{S}_{i}$. This signal is generated via a deterministic function $\psi_{i}:\left(\bigtimes_{j<i}A_{j}\right)\to\mathcal{S}_{i}$ called an experiment. [^32] Abusing notation, we also sometimes take $\psi_{i}:\mathcal{V}^{i-1}\to\mathcal{S}_{i}$, i.e., the experiment maps values to signals instead of actions. For notational convenience, we assume that $\psi_{i}$ is surjective for all $i$. We can think of the experiment as a garbling of the previous bidders’ actions—the experiment can pool together multiple actions from previous bidders to a single signal and so a signal is not always perfectly informative of previous actions. We can recover the public setting with a fully informative experiment, i.e., let $\psi=\mathrm{Id}$, the identity mapping. We can capture classical static game settings via an uninformative experiment that always return the same output, $\psi=\emptyset$. We use $\psi^{-1}_{i}(s_{i})$ to denote the set of $v_{-i}$ that are possible from the perspective of bidder $i$ given its signal.
A revelation principle holds in this setting: for any single-action auction, we can define the direct allocation and transfer rules as $\tilde{x}(v)$ and $\tilde{t}(v)$, respectively, with the appropriate incentive compatibility and individual rationality constraints for real bidders (Lemma D.1 in the Appendix), and appropriate IC constraints for weak and strong shill-proofness (Lemmatta D.2 and D.3, respectively, in the Appendix).
5.2 Single-Action, Ex-Post Auctions Cannot Be Shill-Proof
To finish defining all the terms necessary for our main result of this section, we weaken our notion of ex-post incentive compatiblility in the single-action auction setting to ex-post incentive compatiblility for at least a single bidder: [^33] We exclude the last bidder who takes an action from our definition because a take-it-or-leave-it offer to that bidder can be optimal and ex-post incentive compatible. Theorem 5.2 would still hold if we instead defined mild ex-post incentive compatiblility to mean ex-post incentive compatible for at least two bidders.
Definition 5.1.
A single-action auction is mildly ex-post incentive compatible if for the associated direct mechanism, there exists a real bidder $i<N$ such that truthfulness is an ex-post strategy conditional on the realization of her signal: there exists bidder $i<N$, such that for all $i,v_{i},v_{i}^{\prime},s_{i}$ and $v_{-i},v_{-i}^{\prime}\in\psi_{i}^{-1}(s_{i})$, $\tilde{x}_{i}(v)\cdot v_{i}-\tilde{t}_{i}(v)\geq\tilde{x}_{i}\left(v^{\prime}\right)\cdot v_{i}-\tilde{t}_{i}\left(v^{\prime}\right).$
Theorem 5.2.
@@ 5.2 Single-Action, Ex-Post Auctions Cannot Be Shill-Proof
There exists no single-action, optimal auction that is mildly ex-post incentive compatible and weakly shill-proof.
Proof Sketch.
Consider any real bidder $i<N$. By weak shill-proofness, the transfer from bidder $i$, conditional on winning (or losing) the auction, is invariant to the values of bidders who take actions after her (Lemma D.4 in the Appendix). If this were not the case, then if every bidder $j>i$ is a shill bidder, the shill bidders would report the values that would maximize the transfer from the winning bidder. By mild ex-post incentive compatiblility, the transfer from bidder $i$, conditional on winning the auction, is invariant to her value (Lemma D.5 in the Appendix). This is because if there were multiple winning reports with different transfer amounts, only the smallest transfer amount would make truthful reporting of the value an ex-post strategy. So, in every single-action, optimal auction, the transfer from the winning bidder $i$, can depend only on the values reported by bidders before $i$. But, this means that if a bidder has positive utility for winning the item (as would be the case if $v_{i}>v_{j}$ for all $j<i$), then she should report $v^{M}$ to maximize the probability of winning (without changing the transfer paid upon winning). Thus, the auction must treat bidder $i$ as if she reported $v^{M}$, which violates the allocation rule of an optimal auction. ∎
@@ 6 Discussion
As discussed in the review of literature, another notion of “cheating” by the auctioneer is that of in-credibility introduced by Akbarpour and Li (2020). An auction is credible if a revenue-maximizing auctioneer has no incentive to lie about what other players are doing. That information environment differs from this paper because we assume that bidders correctly (though perhaps not fully) perceive the actions of other players and where in the game tree they are. In the Online Appendix, we formally define credibility in our setting (Definition O.6) and prove the following implications:
@@
Suppose $(G,\sigma)$ is an optimal auction. If the auction is strongly shill-proof, then it must be credible. If the auction is credible, then it must be weakly shill-proof.
We also define $\psi$-credibility for single-action auctions (Definition O.8) as a generalization of credibility allowing for bidders to have exogenous signals (where $\psi$ is our notation from Section 5) about the actions of other bidders as well as additional communication from the auctioneer. In a single-action auction, we define $\psi=\mathrm{Id}$ to mean that $\psi_{i}(a_{j<i})=a_{j<i}$, i.e., that the signals reveal the actions of previous bidders. We define $\psi=\emptyset$ to mean the opposite, $\psi_{i}(a_{j<i})=\emptyset$. This is the classic static auction setting. We prove that the implications of Proposition 6.1 still hold and give conditions under which credibility coincides with strong and weak shill-proofness:
@@
Suppose $(G,\sigma)$ is a single-action, optimal auction. If the auction is strongly shill-proof, then it must be $\psi$-credible. If it is $(\psi=\emptyset)$-credible, then it is strongly shill-proof. If the auction is $\psi$-credible, then it must be weakly shill-proof. If it is weakly shill-proof, then it is $(\psi=\mathrm{Id})$-credible.
Proposition 6.2 implies that Theorem 5.2 is a generalization of the credible trilemma (Akbarpour and Li, 2020, Theorem 1).

### v3

@@ 4.2.2 Weakly Shill-Proof and Strategy-Proof Auctions
Proposition 4.12.
For all $\varepsilon>0$, there exists a value distribution $F$ and screen level $\theta^{Y}$ such that $Q^{AS,Y}(F)/Q^{E}(F)<\varepsilon$ and the ascending, screening auction with screening level $\theta^{Y}$ is weakly shill-proof, ex-post incentive compatible, and optimal.
The ascending, screening auction is orderly and optimal by construction; and it is ex-post incentive compatible because the English auction phase and the second-price phase both induce the same (ex-post incentive compatible) allocation and transfer rule. The larger $\theta^{Y}$ is, the less that can be extracted in expectation from shill bidding and the more likely it is that a shill bidder will win the item if she shill bids. We provide a sufficient minimum bound on $\theta^{Y}$ based on a few moments of a distribution (not its number of atoms), such that for distributions with “thin-enough” right tails—in particular monotone hazard rate distributions—the ascending, screening auction is weakly shill-proof (Lemma C.10). We can then construct a sequence of distributions with increasing numbers of atoms and constant $\theta^{Y}$ to complete the proof. The distributional requirements we need here differ from the class we find to prove Proposition 4.6, demonstrating that designing weakly shill-proof auctions has rich interactions with properties of the prior distribution.
5 Affiliation, Interdependent Values, and Shill-Proofness
In this section, we generalize our model to allow for interdependent values and a more general distribution of shill bidders. We will then show that shill-proofness admits an order with respect to the type distribution and value function: the higher the level of affiliation and level of common values that is present in the environment, the harder it is for an auction format to be shill-proof, either weakly or strongly. Next, we show that Lemma 3.1 holds under our general model and pins down the optimal allocation rule for strongly shill-proof auctions in affiliated environments. Finally, we show that our main results (Theorems 3.6 and 4.3) still hold under our more general model.
5.1 Generalized Model
We maintain the assumptions that there is a set of potential bidders $B$, with $\left|B\right|=N$, who might participate and that the seller’s value is commonly known to be 0. We also maintain our previous assumptions on real and shill bidder utilities and information sets. However, instead of assuming that each potential bidder has probability $p$ of participating, we instead assume that the set of real bidders is drawn symmetrically and randomly, i.e., for any $R^{\prime},R^{\prime\prime}\subset B$, if $|R^{\prime}|=|R^{\prime\prime}|$, then $\mathds{P}\left[R=R^{\prime}\right]=\mathds{P}\left[R=R^{\prime\prime}\right]$. For parsimony, we also assume full support, but the probability of certain sets of real bidders can be arbitrarily small; this rules out mechanisms that condition on large coalition of bidders in unusual ways in order to become weakly shill-proof. [^30] Our results for strong shill-proofness do not require a full support distribution over shill bidder sets. Letting the probability of certain realizations go to zero (or relaxing the full support assumption entirely), this treatment of the set of shill bidders nests many natural models of shill bidders. For example, it nests a model of shill bidding where there is at most or exactly one shill bidder. It also nests our model of shill bidding from Section 2 where each bidder has an i.i.d. chance of being a real bidder.
In our general model, we interpret the type $\theta_{i}\in\vartheta$ to flexibly represent both a private value for the item and a signal of the common value for the item. We will expand on exactly how the type enters the valuation for real bidders below. As notation, for each shill bidder $i\in S$, we assign their type to be $\theta_{i}=0$ to represent them having the lowest possible valuation for the item. Then, we can say types are jointly drawn from a full-support, symmetric $F$ and we define $f(\theta)=\mathds{P}_{w\sim F}\left[w=\theta\right]$ to be the pmf of the distribution. We assume that the type distribution is affiliated in the Milgrom and Weber (1982) sense: For all $\theta,\theta^{\prime}\in\vartheta^{N}$,
$\log f(\theta\vee\theta^{\prime})+\log f(\theta\wedge\theta^{\prime})\geq\log f(\theta)+\log f(\theta^{\prime}).$
Affiliation captures the idea that types are positively correlated: the probability of types/signals all being high or low is higher than the probability of some signals being high and some being low.
There is a value function $v:\vartheta\times\vartheta^{N-1}\to\mathbb{R}$ such that for each bidder $i$, her value for the item is $v\left(\theta_{i},\left\{\theta_{j}\right\}_{j\neq i}\right)$; note that this form implies that bidders have symmetric preferences over the types of other bidders. We also make a few functional form restrictions on $v$ below.
Assumption 5.1.
The value function $v$ for a bidder satisfies the following properties:
  1. (i)
    For all $k$, $v(\theta^{k},0)=\theta^{k}$;
  2. (ii)
    $v$ is weakly increasing in $\theta_{-i}$;
  3. (iii)
    If $\theta_{i}\geq\theta_{j}$, then $v\left(\theta_{i},\left\{\theta_{j},\theta_{-(i,j)}\right\}\right)\geq v\left(\theta_{j},\left\{\theta_{i},\theta_{-(i,j)}\right\}\right)$; and
  4. (iv)
    For all $\theta$ such that $\theta_{i}\geq\max_{j\neq i}\left\{\theta_{j}\right\}$, $v$ is weakly super-modular and has weakly decreasing differences.
Condition (i) is a (without loss) normalization conditional on the other parts of the assumption. Condition (ii) means that $\theta$ encodes common value preferences—a bidder values an item more when other bidders value it more. Condition (iii) means that bidders value their own high-type realizations more than they value high-type realizations for other bidders. (This is the standard single-crossing assumption needed for a responsive, incentive compatible mechanism to exist.) Condition (iv) means that conditional on having the highest ex-ante signal of value, a bidder has higher value for other bidders’ signals whenever her type is high and has diminishing marginal returns to high signals.
Note that the standard, private values setting where for all $\theta$, $v\left(\theta_{i},\theta_{-i}\right)=\theta_{i}$ satisfies Assumption 5.1. Most other value functions in the literature also satisfy these conditions; for example, generalized, additive, interdependent values $v(\theta)=\theta_{i}+\kappa\sum_{j\neq i}\theta_{j}$ with $\kappa\leq 1$ and maximum common values $v(\theta)=\max_{i}\left\{\theta_{i}\right\}$.
5.2 Shill-Proof Order
To prove that increasing the level of affiliation and interdependence of values makes it more difficult for an auction to be shill-proof, let us begin by formally defining what it means for the value distribution to be “more affiliated” and for the value function to be “more commonly valued.”
Definition 5.2.
Consider two distributions $F,F^{\prime}\in\Delta(\vartheta^{N})$ such that all marginal distributions are the same: $F_{i}=F^{\prime}_{i}$, for all $i$. A distribution $F^{\prime}$ is more affiliated than $F$, $F^{\prime}\succeq_{\mathrm{Aff}}F$, if for all $x,y\in\vartheta^{N}$, $\log f^{\prime}(x\vee y)-\log f^{\prime}(x)\geq\log f(y)-\log f(x\wedge y)$.
Our definition of the affiliation order is standard and from Karlin and Rinott (1980). When $F^{\prime}\succeq_{\mathrm{Aff}}F$, values between bidders are more highly correlated under $F^{\prime}$ than $F$. We fix the marginal distributions to emphasize that our definition focuses on the correlation between bidders and not the relative strength of the distributions. All affiliated distributions are more affiliated than the type distribution where bidders are independent: If $F$ is affiliated, then $F\succeq_{\mathrm{Aff}}\prod_{i}F_{i}$.
Definition 5.3.
A value function $v^{\prime}$ is more commonly valued than $v$, $v^{\prime}\succeq_{\mathrm{Com}}v$, if $v^{\prime}$ is weakly more super-modular than $v$ and for all $\theta$, $v^{\prime}(\theta)\geq v(\theta)$.
The point-wise comparison in Definition 5.3 makes sense because $v,v^{\prime}$ are normalized so that $v(\theta^{k},0)=v^{\prime}(\theta^{k},0)=\theta^{k}$—and therefore the point-wise comparison means that other bidders’ signals are valued more as the value function becomes more commonly valued. The increase in super-modularity means that we also require types to be more complementary when we say that the value function is more commonly valued. We view this as reasonable because if signals became less complementary, there would exist $v^{\prime}\succeq_{\mathrm{Com}}v$ such the marginal change in bidder $i$’s value based on her own type would be less under $v^{\prime}$ than under $v$. Any value function $v$ satisfying Assumption 5.1 is more commonly valued than the value function $v_{\mathrm{private}}(\theta_{i},\theta_{-i})=\theta_{i}$ arising under the private values model.
Letting $\tilde{t}^{*}(\tilde{x},\mu,\xi_{0},v,F)$ be defined as some optimal transfer rule conditional on the extensive-form game and the primitives of the environment, we are now in position to formally state our main result for this section:
@@ 5.2 Shill-Proof Order
Consider affiliated type distributions $F$ and $F^{\prime}$, and value functions $v$ and $v^{\prime}$ satisfying Assumption 5.1. Suppose $(\tilde{x},\tilde{t}^{*}(\tilde{x},\mu,\xi_{0},v,F),\mu,\xi_{0})$ is orderly and strongly shill-proof. Then, if $v^{\prime}\succeq_{\mathrm{Com}}v$ and $F\succeq_{\mathrm{Aff}}F^{\prime}$, it is the case that $(\tilde{x},\tilde{t}^{*}(\tilde{x},\mu,\xi_{0},v^{\prime},F^{\prime}),\mu,\xi_{0})$ is strongly shill-proof. The same statement holds for weak shill-proofness.
Proof Sketch..
To begin the proof sketch, let us first observe that the ex-interim transfers from each bidder is pinned down based on the ex-interim allocation rule via a bidder’s incentive compatibility constraint.
Lemma 5.5.
For every auction $(\tilde{x},\tilde{t}^{*}(\tilde{x},\mu,\xi_{0},v,F),\mu,\xi_{0})$, the ex-interim transfer rule for bidder $i$ is
$T_{i}(\theta_{i};\Theta)=X_{i}(\theta_{i};\Theta)v^{*}(\theta_{i},\theta_{i})-\sum_{m:\theta^{j_{m}}<\theta_{i}}X_{i}(\theta^{j_{m}};\Theta)\cdot(v^{*}(\theta^{j_{m}},\theta^{j_{m+1}};\Theta)-v^{*}(\theta^{j_{m}},\theta^{j_{m}};\Theta)),$
 $\text{ where }v^{*}_{i}(\theta^{k},\theta^{k^{\prime}};\Theta)=\mathds{E}\left[v(\theta^{k},\theta_{-i})\mid\theta_{i}=\theta^{k^{\prime}},\theta_{-i}\in\Theta_{-i},\tilde{x}_{i}(\theta^{k},\theta_{-i})=1\right]$
is the expected value of the item for bidder ii of type $\theta^{k^{\prime}}$ conditional on winning the item and playing as if she were type $\theta^{k}$ and $\left\{\theta^{j_{m}}\right\}_{m}$ are the ordered atoms of $\Theta_{i}$.
Note that feasibility normally pins down the transfer rule via the envelope theorem. However, we are working with discrete types instead of continuous types and so the incentive compatibility constraint is slackened as the change in allocation probability from misreporting is discrete. So, we use optimality to select among this multitude of ex-interim transfers that maintain incentive compatibility. (This is an adaptation of the standard envelope theorem formulation of ex-interim transfers adapted to the extensive-form game by noting that if, at any point in the game, $T_{i}$ does not take this form, there exists a profitable deviation by bidder $i$ to commit to acting as a lower type throughout the rest of the game.)
Next, we prove the following technical lemma.
Lemma 5.6.
Consider sets $S$ and $S^{\prime}$ such that $S\cap S^{\prime}=\emptyset$ and $\min_{v\in S}\left\{v\right\}>\min_{v\in S^{\prime}}\left\{v\right\}$. Then, for any weakly super-modular, non-decreasing function $g$ with decreasing differences, if $F^{\prime}\succeq_{\mathrm{Aff}}F$, then for all $i$,
$\mathds{E}_{\theta\sim F^{\prime}}\left[g(\theta)\mid\theta_{i}\in S\right]-\mathds{E}_{\theta\sim F^{\prime}}\left[g(\theta)\mid\theta_{i}\in S^{\prime}\right]\geq\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S\right]-\mathds{E}_{\theta\sim F}\left[g(\theta)\mid\theta_{i}\in S^{\prime}\right].$ (1)
The proof then proceeds by contrapositive for affiliation and then common values. We consider that the auction is not weakly shill-proof for distribution $F$ and consider $F^{\prime}\succeq_{\mathrm{Aff}}F$. We show that if there exists a profitable deviation under $F$, then that same deviation must be profitable under $F^{\prime}$. Any profitable deviation either manipulates a bidder’s ex-interim expected transfer or ex-post manipulates her payment. Given that we have fixed the extensive form, any deviation that is possible under $(v,F)$ is possible under $(v^{\prime},F^{\prime})$. So, ex-post manipulation will be profitable under both and so we only need to consider ex-interim manipulations. We can re-write expected profit of deviation as the difference between the expectation of a sum taken over the true and manipulated type reports. We show that the sum satisfies the conditions of Lemma 5.6 and so we can conclude that deviating increases profits. Note that the probability that bidder $i$ wins the item is constant between the two type distributions because the allocation rule is held constant and so we have shown that the auction is not weakly shill-proof under $F^{\prime}$. The argument when strongly shill-proof is the exact same. The proof when considering common values is also by contrapositive. In this case, we apply the fact that $v^{\prime}$ is more super-modular and point-wise larger to directly conclude that the profit from the deviation is larger. ∎
5.3 Extending our Characterizations
We can now use Theorem 5.4 to generalize our main characterization results from previous sections. To begin, we observe that while we know the optimal allocation rule in an IPV environment and the revenue equivalence theorem implies that any transfer rule yields the same revenue, the same need not hold with interdependent values and affiliation. So, we must simultaneously search for an allocation and transfer rule that satisfies our desiderata.
In order to find the optimal allocation and transfer rules, we first pin down the transfer rule and then the allocation rule. Our proof of Lemma 3.1 holds in our general environment and so strong shill-proofness implies that the transfer rule must be pay-as-bid. Then, having pinned down the transfer rule, we can see first that the unconstrained efficient allocation is feasible, and also that in the optimal auction we will never want to allocate to a bidder who does not have the highest type. Thus, we can solve the monopolist screening problem to find the optimal reserve in this auction. The formal result is as follows.
Proposition 5.7.
Under any affiliated distribution and value function satisfying Assumption 5.1, if an auction $(\tilde{x},\tilde{t},\mu,\xi_{0})$ is strongly shill-proof, then it is pay-as-bid: For all $\xi$, $v_{\xi}$, $\theta_{-\xi}$, and $\theta_{-\xi}^{\prime}$,
$\tilde{x}_{\xi}\left(\theta_{\xi},\theta_{-\xi}\right)=\tilde{x}_{\xi}\left(\theta_{\xi},\theta_{-\xi}^{\prime}\right)\implies\tilde{t}_{\xi}\left(\theta_{\xi},\theta_{-\xi}\right)=\tilde{t}_{\xi}\left(\theta_{\xi},\theta_{-\xi}^{\prime}\right).$
All efficient mechanisms have an allocation rule such that
$\sum_{i}\tilde{x}^{E}_{i}(\theta)=1\text{ and }i\notin\operatorname*{argmax}_{j}\left\{\theta_{j}\right\}\implies\tilde{x}^{E}_{i}(\theta)=0.$
All optimal mechanisms have an allocation rule such that
$\sum_{i}\tilde{x}^{*}_{i}(\theta)=\begin{cases}1&\max_{j}\left\{\theta_{j}\right\}\geq\theta^{\rho^{*}}\\ 0&\text{otherwise}\end{cases}\text{ and }i\notin\operatorname*{argmax}_{j}\left\{\theta_{j}\right\}\implies\tilde{x}^{E}_{i}(\theta)=0$
for some $\theta^{\rho^{*}}$.
We can now generalize Theorem 3.6. Modifying the transfer rule in the Dutch auction to be the appropriate pay-as-bid rule, the following corollary holds.
Corollary 5.8.
Under any affiliated type distribution and value function satisfying Assumption 5.1, the statement of Theorem 3.6 holds.
Proof.
The proof that any Dutch auction is strongly shill-proof is the exact same as in the IPV case. By Proposition 5.7, we know what the orderly and optimal allocation rule is. By Theorem 3.6, we know that in the IPV case, no non-Dutch auction will be strongly shill-proof. We then apply Theorem 5.4 to conclude that all non-Dutch auctions are not strongly shill-proof and optimal in general. ∎
Many of our results for weakly shill-proof auctions continue to hold in our general model when we restrict to optimal transfer rules. The generalization of Theorem 4.3 is as follows.
Corollary 5.9.
@@ 5.3 Extending our Characterizations
Under any affiliated type distribution $F$ and value function $v$ satisfying Assumption 5.1, there exists no orderly auction $(\tilde{x},\tilde{t}^{*}(\tilde{x},\mu,\xi_{0},v,F),\mu,\xi_{0})$ that is single-action, mildly ex-post incentive compatible and weakly shill-proof.
Proof.
Applying Theorem 5.4 to Theorem 4.3 with a restriction on possible transfer rules completes the proof. ∎
While regularity is not as applicable with affiliation, our robustness result (Corollary 4.8) for the Dutch auction still hold.
Corollary 5.10.
Under any affiliated type distribution $F$ and value function $v$ satisfying Assumption 5.1 and for any public and efficient auction $(\tilde{x},\tilde{t}^{*}(\tilde{x},\mu,\xi_{0},v,F),\mu,\xi_{0})$, if the auction is not a Dutch auction with reserve price 0, there exists a value distribution under which the auction is not weakly shill-proof.
Proof.
Since the Dutch auction with reserve price 0 is strongly shill-proof, it must be weakly shill-proof. We then apply Theorem 5.4 to Corollary 4.8 to rule out any other auction format being robustly weakly shill-proof and complete the proof. ∎
We conclude this section by observing that Proposition 4.12 does not necessarily hold in the general model. In fact, as we mentioned in the introduction, even the English auction is not necessarily weakly shill-proof under an IPV type distribution if the distribution is not regular. We leave as a future research question to find conditions on type distributions such that some strategy-proof and weakly shill-proof mechanism exists.
@@ 6 Discussion
As discussed in the review of literature, another notion of “cheating” by the auctioneer is that of in-credibility introduced by Akbarpour and Li (2020). An auction is credible if a revenue-maximizing auctioneer has no incentive to lie about what other players are doing. That information environment differs from this paper because we assume that bidders correctly (though perhaps not fully) perceive the actions of other players and where in the game tree they are. In the Online Appendix, we formally define credibility in our setting (Definition E.8) and prove the following implications:
@@
Suppose $(G,\sigma)$ is an orderly auction. If the auction is strongly shill-proof, then it must be credible. If the auction is credible, then it must be weakly shill-proof.
We also define $\psi$-credibility for single-action auctions (Definition E.10) as a generalization of credibility allowing for bidders to have exogenous signals (where $\psi$ is our notation from Section 4.1) about the actions of other bidders as well as additional communication from the auctioneer. Recall that we defined $\psi=\mathrm{Id}$ to mean that the signals reveal the actions of previous bidders and $\psi=\emptyset$ to be static auction setting. We prove that the implications of Proposition 6.1 still hold and give conditions under which credibility coincides with strong and weak shill-proofness:
@@
Suppose $(G,\sigma)$ is a single-action, orderly auction. If the auction is strongly shill-proof, then it must be $\psi$-credible. If it is $(\psi=\emptyset)$-credible, then it is strongly shill-proof. If the auction is $\psi$-credible, then it must be weakly shill-proof. If it is weakly shill-proof, then it is $(\psi=\mathrm{Id})$-credible.
Proposition 6.2 implies that Theorem 4.3 is a generalization of the credibility trilemma (Akbarpour and Li, 2020, Theorem 1).